# CCSBench: Evaluating Compositional Controllability in LLMs for Scientific Document Summarization

### v1

@@ Abstract
To broaden the dissemination of scientific knowledge to diverse audiences, scientific document summarization must simultaneously control multiple attributes such as length and empirical focus. However, existing research typically focuses on controlling single attributes, leaving the compositional control of multiple attributes underexplored. To address this gap, we introduce CCSBench, a benchmark for compositional controllable summarization in the scientific domain. Our benchmark enables fine-grained control over both explicit attributes (e.g., length), which are objective and straightforward, and implicit attributes (e.g., empirical focus), which are more subjective and conceptual. We conduct extensive experiments on GPT-4, LLaMA2, and other popular LLMs under various settings. Our findings reveal significant limitations in large language models’ ability to balance trade-offs between control attributes, especially implicit ones that require deeper understanding and abstract reasoning.  [^1] Our data is available at: https://huggingface.co/datasets/dyxohjl666/CCSBench.
CCSBench: Evaluating Compositional Controllability in LLMs for Scientific Document Summarization
Yixi Ding1    Jiaying Wu1thanks: Corresponding authors.  Tongyao Zhu1  Yanxia Qin2  Qian Liu3  Min-Yen Kan [^1] footnotemark: 1 1National University of Singapore 2Singapore University of Technology and Design 3 Sea AI Lab {yixi.d,kanmy}@comp.nus.edu.sgjiayingw@nus.edu.sgtongyao.zhu@u.nus.edu yanxia_qin@sutd.edu.sgliuqian@sea.com
1 Introduction
Compositional controllability – the ability to generate tailored summaries based on multiple attributes – is essential for scientific document summarization. For instance, different audiences require different summaries of a scientific paper on a newly proposed language model. As illustrated in Figure 1, an NLP beginner might prefer an understandable, concept-focused explanation, while a product manager might prioritize a brief overview of the model’s empirical performance. The flexibility of compositional control enhances user satisfaction by tailoring summaries to specific requirements across diverse expertise levels and interests.
@@ 1 Introduction
Figure 1: Compositional controllability allows scientific summaries to be tailored to diverse reader needs, addressing specific reader requirements.
@@
Despite the widespread need for compositional controllability in scientific summarization, this desirable feature remains underexplored in controllable text generation research. Most current efforts focus on using large language models (LLMs) to generate non-academic texts such as movie reviews Liu et al. (2024) and news articles Chen and Shu (2024), which are less complex than scientific documents adhering to strict academic conventions. Preliminary investigations into compositional scientific summarization Ding et al. (2023) address only straightforward control requirements such as length and keyword inclusion, overlooking more subtle user requirements such as focus on empirical aspects. To develop methods suitable for real-world deployment, it is non-trivial to create benchmarks that capture more complex control attributes and reflect them jointly in a compositional manner.
@@
In this paper, we introduce CCSBench, a benchmark for compositional scientific summarization that enables control over diverse fine-grained attributes. Inspired by Kahneman’s cognitive theory Kahneman (2013), which describes human cognition operating in two modes—the fast, intuitive “System 1” and the slow, deliberate “System 2”, CCSBench incorporates both explicit control attributes that are objective and straightforward, and implicit control attributes that are conceptual and require deeper reasoning. Explicit attributes, such as length, are easily quantifiable and straightforward for both language models and humans to process. In contrast, implicit attributes, such as empirical focus, are subjective and necessitate more human-like understanding and reasoning.
@@
As illustrated in Figure 2, CCSBench is structured around two explicit attributes: (1) length and (2) keyword inclusion, and two implicit attributes: (3) readability and (4) focus. The length attribute sets the target word count for the summary, while keyword inclusion ensures the presence of specific terms in the output. The readability attribute controls the complexity and accessibility of the language, tailoring the text to the comprehension level of a specific audience, whether technical or general. The focus attribute, short for “Empirical Focus Level”, instructs the model to prioritize empirical evidence over conceptual aspects, necessitating deeper contextual understanding and abstract reasoning. By adjusting combinations of these attributes, we curate diverse scientific summaries – such as concise, highly readable conceptual explanations – that meet real-world needs and cater to diverse reader expectations.
@@
We evaluate a diverse set of representative LLMs on CCSBench, including black-box models and smaller fine-tuned variants. The results reveal significant limitations in LLMs’ ability to manage trade-offs between control attributes, especially due to inadequate abstract reasoning when handling implicit attributes. In parameter-efficient fine-tuning (PEFT) settings, we find that decoder-only models like LLaMA2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) struggle due to their inability to model long-term dependencies, while encoder-decoder models like Flan-T5 (Chung et al., 2022) adapt better to the task. Further analysis shows that popular task composition methods, such as AdapterFusion Pfeiffer et al. (2021), are unsuitable for our task due to oversimplified aggregation among adapters. Our findings suggest the need for more targeted research on compositionality in scientific summarization.
Figure 2: Illustration of explicit and implicit control attributes in CCSBench.
2 Related Work
2.1 Controllable Summarization
@@
Controllable summarization has garnered substantial interest, with models targeting aspects such as length, topic, keywords, and sentiment. Early research primarily examined the influence of oracle attributes on summary quality Fan et al. (2018); See et al. (2017). More recent studies have focused on controlling individual attributes such as length and keywords Chan et al. (2021); He et al. (2020), but these approaches typically optimize one attribute at a time, limiting their effectiveness when handling multiple attributes simultaneously. MacSum Zhang et al. (2023) investigates multi-attribute control in news and dialogue domains, where the content is less complex than in scientific texts. In the scientific domain, compositional controllable summarization presents distinct challenges due to the technical complexity and granularity of the language. CocoScisum Ding et al. (2023) attempted to control both length and keywords in scientific papers but was limited to these two explicit attributes. Our work, CCSBench, addresses this limitation by providing fine-grained control over both explicit and implicit attributes, enabling a more comprehensive evaluation of scientific summaries.
2.2 LLM Instruction Following
LLMs demonstrate impressive capabilities in following instructions and adapting to various downstream tasks through prompt-based inputs Zhao et al. (2024). TThis flexibility allows LLMs to generate controlled outputs with minimal supervision, making them suitable for tasks requiring specific constraints or user-driven outcomes. Recent studies have investigated LLMs’ effectiveness in controllable text generation, particularly for single-attribute tasks. For instance, research shows that LLMs can control attributes like length Li et al. (2024) and formality Zhou et al. (2024), enabling precise adjustments based on user instructions. These findings underscore LLMs’ adaptability in responding to control signals in general text generation. However, despite the critical need for concise, readable, and targeted summaries in scientific domains, the challenge of simultaneously managing multiple attributes remains largely unaddressed. In this work, we provide a comprehensive evaluation of LLMs’ compositional controllability for scientific summarization.
3 CCSBench Data Curation
3.1 Task Formulation
@@
Length
specifies the desired number of words in the summary, ranging from concise to detailed, and catering to different audience needs based on their familiarity with the topic or time constraints. Length is divided into five bins, each representing a 50-word range (e.g., Bin 0: 0-50 words; Bin 1: 51-100 words).
Keywords
ensures the inclusion of key terms from the source document, enabling readers to quickly grasp the core elements and assess the paper’s relevance. Keywords can be left empty or set as pre-defined terms to be included in the summary.
Readability
controls the syntactic complexity and vocabulary to match the audience’s proficiency. Controlling readability allows summaries to be either more accessible or technical, depending on the audience, which enhances comprehension and engagement. Readability can take two values: [normal, high].
Empirical Focus Level (Focus)
adjusts the emphasis on empirical results versus theoretical and conceptual contributions. Focus can take two values: [low, high]. A high empirical focus highlights practical, data-driven aspects, such as data collection and experimental results. A low empirical focus emphasizes theoretical contributions, such as frameworks, hypotheses, and broader implications. This control allows the summary to meet different reader objectives, whether they seek practical outcomes or conceptual insights.
3.2 Dataset
Our dataset  CCSBench is built on top of the arXiv dataset Clement et al. (2019), sourced from the widely used online pre-print repository, arXiv.  CCSBench comprises scientific papers from the subjects of Computer Science or Artificial Intelligence. To ensure high-quality input with minimal noise, we use the introduction and conclusion sections (I+C, Meng et al., 2021) of each document as the input text. The benchmark consists of four single-attribute control datasets and one compositional control dataset, with approximately 4,000 instances in each dataset.
We divide the dataset into training, validation, and test sets using a 60/20/20 split. The statistics for each dataset are presented in Table 1.
@@ 3.2 Dataset
Table 1: CCSBench statistics.
3.3 Dataset Construction
@@
We begin with KWX Fu (2023), a dataset that provides keywords for each scientific document sourced from arXiv. From this, we select 3,700 papers in the subjects of Computer Science and Artificial Intelligence to form the foundation of our controllable summarization dataset, using their abstracts as natural reference summaries.
@@
Dataset FAC FLU CTRL ROUGE-1 ROUGE-2 ROUGE-L
Readability 0.98 0.93 0.92 53.88 21.81 41.31
Focus 0.91 0.93 0.93 51.25 34.23 41.41
Table 2: Evaluation results validate the high quality of LLM-generated summaries in CCSBench.
@@
Validation of Data Quality
We validate the quality of the GPT-4-generated summaries in CCSBench through both automated and human evaluations. First, we calculate the ROUGE scores between the generated and original summaries to assess fluency and content accuracy. Next, we perform a human evaluation to further verify the controllability and overall quality of the generated data.
We sample 120 documents each from $D_{read}$ and $D_{focus}$, and employ human workers from Amazon Mechanical Turk [^2] https://www.mturk.com/ for binary assessments on three key aspects: (1) Factual Consistency (FAC): whether the generated summary accurately reflects the key points of the reference, (2) Fluency (FLU): whether the generated summary is written in natural, fluent language, and (3) Controllability (CTRL): for Readability, whether the generated summary is easier to read than the reference; and for Focus, whether empirically-focused summaries emphasize data collection, experimental setup, and research results more than the conceptually-focused summaries (detailed descriptions are provided in Appendix C). The evaluation results, shown in Table 2, demonstrate high accuracies exceeding 0.9 across all metrics, which indicate the high quality of generated data in CCSBench.
4 Experimental Setup
4.1 Models
We evaluate a range of popular models, both closed-source and open-source, including GPT-4 (OpenAI et al., 2024), LLaMA2  (Touvron et al., 2023), Mistral (Jiang et al., 2023), and Flan-T5 (Chung et al., 2022), to assess their capabilities in compositional controllable summarization. For open-source models, we also experiment with parameter-efficient fine-tuning (PEFT) using Low-Rank Adaptation (LoRA) Hu et al. (2021) to adapt them to our dataset. Additionally, we revisit baselines methods for controllable summarization from the news domain, including hard prompt (HP) Zhang et al. (2023) and soft prefix tuning (SP) Li and Liang (2021). Further implementation details, including training setups, are provided in Appendix A.
4.2 Evaluation Metrics
We evaluate all models based on both summarization quality and attribute controllability. We employ ROUGE Lin (2004) to evaluate the overall quality. To assess attribute controllability, we employ the following metrics tailored to each attribute.
@@ 4.2 Evaluation Metrics
We measure length controllability using the Mean Absolute Deviation (MAD,  Liu et al., 2018) and the Pearson Correlation Coefficient (PCC,  Liu et al., 2018) to evaluate the distance between the target and generated lengths.
@@
Keyword control is evaluated using the Success Rate (SR, He et al., 2020), which measures the fraction of specified keywords present in the generated summaries through exact matching after stemming.
@@
For readability, we calculate the Flesch-Kincaid Grade Level (FKGL,  Kincaid et al., 1975), where a lower score indicates higher readability. The difference in FKGL scores ($\delta_{FKGL}$) between the high readability setting ($F_{high}$) and the normal readability setting ($F_{normal}$) reflects the model’s ability to control this attribute. A higher $\delta_{FKGL}$ is desirable.
@@
For empirical focus, we use GPT-4 to classify summaries into high or low empirical focus levels. Validated with 93% accuracy on a manually labeled dataset, GPT-4 predictions are used to compute the F1 score for each category, assessing the model’s focus control. Prompt details are provided in Appendix B.
@@
Table 3: Performance comparison of language models on CCSBench, evaluating overall quality and four control attributes. Arch refers to model architectures, with ED for encoder-decoder and DO for decoder-only. Bold numbers indicate the best performance, while underlined numbers represent the second-best.
5 Experiments

### v2

@@ Abstract.
To broaden the dissemination of scientific knowledge to diverse audiences, it is desirable for scientific document summarization systems to simultaneously control multiple attributes such as length and empirical focus. However, existing research typically focuses on controlling single attributes, leaving the compositional control of multiple attributes underexplored. To address this gap, we introduce CCSBench, the first evaluation benchmark for compositional controllable summarization in the scientific domain. Our benchmark enables fine-grained control over both explicit attributes (e.g., length), which are objective and straightforward, and implicit attributes (e.g., conceptual or empirical focus), which are more subjective and abstract. We conduct extensive experiments using various large language models (LLMs) under various settings, including in-context learning, parameter-efficient fine-tuning, and two-stage modular methods for balancing control over different attributes. Our findings reveal significant limitations in LLMs capabilities in balancing trade-offs between control attributes, especially implicit ones that require deeper understanding and abstract reasoning.  [^1] Data is available at: https://huggingface.co/datasets/dyxohjl666/CCSBench.
Abstractive Summarization, Scientific Document Summarization, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: KDD’25 SciSoc LLM Workshop: Large Language Models for Scientific and Societal Advances; August 6, 2025; Toronto, ON, Canadaccs: Computing methodologies Natural language processingccs: Information systems Summarization
1. Introduction
Compositional controllability – the ability to generate tailored summaries based on multiple attributes – is essential for scientific document summarization systems to effectively communicate research to diverse audiences. Scientific papers contain complex information, but different readers have distinct needs and preferences when engaging with this content. For instance, as illustrated in Figure 1, an NLP beginner may seek a clear, concept-focused explanation, while a product manager may prioritize a concise summary highlighting the model’s empirical performance. By dynamically adjusting summaries to align with varying expertise levels and interests, compositional control ensures that scientific knowledge is more accessible, relevant, and engaging for a broader audience.
@@ 1. Introduction
Figure 1. Compositional controllability allows scientific summaries to be tailored to diverse reader needs, addressing specific reader requirements.
@@
Despite the broad demand for compositional controllability in scientific summarization, this capability remains largely underexplored in controllable text generation research. Most current efforts focus on using large language models (LLMs) to generate non-academic texts such as movie reviews (Liu et al., 2024a) and news articles (Chen and Shu, 2024), which are less complex than scientific documents adhering to strict academic conventions. Preliminary investigations into compositional scientific summarization (Ding et al., 2023) address only straightforward control requirements such as length and keyword inclusion, overlooking more subtle user requirements such as focus on empirical aspects. Developing methods for real-world deployment requires benchmarks that not only incorporate more sophisticated control attributes but also capture their interplay in a truly compositional manner, a challenge that remains unresolved.
@@
In this paper, we introduce CCSBench,the first evaluation benchmark for compositional scientific summarization that enables fine-grained control over multiple attributes. Drawing inspiration from Kahneman’s cognitive theory (Kahneman, 2013), which distinguishes between the fast, intuitive “System 1” and the slow, deliberate “System 2” modes of thinking, CCSBench incorporates both explicit and implicit control attributes. Explicit attributes (e.g., length) are objective, easily quantifiable, and straightforward for both LLMs and humans to process. In contrast, implicit attributes (e.g., empirical focus) require deeper reasoning and human-like understanding, making them more challenging for LLMs to control effectively.
@@
As illustrated in Figure 2, CCSBench is structured around two explicit attributes: (1) length and (2) keyword inclusion, and two implicit attributes: (3) readability and (4) focus. Length sets the target word count for the summary, while keyword inclusion ensures the presence of specific terms in the output. Readability controls language complexity and accessibility, tailoring the summary for technical or general audiences. Focus, short for “Empirical Focus Level”, adjusts the emphasis on empirical evidence versus conceptual aspects, requiring deeper contextual understanding and abstract reasoning. By dynamically adjusting these attributes, we curate diverse scientific summaries – such as concise, highly readable conceptual explanations – that meet real-world needs and cater to diverse reader expectations.
@@
We evaluate a diverse set of LLMs on CCSBench, including both proprietary LLMs and fine-tuned open-source variants. Results reveal significant limitations in managing trade-offs between control attributes, particularly due to inadequate abstract reasoning for implicit attributes. In parameter-efficient fine-tuning (PEFT), decoder-only models like LLaMA2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) struggle with long-term dependencies, while encoder-decoder models like Flan-T5 (Chung et al., 2022) adapt more effectively. Further analysis shows that standard task composition methods, such as AdapterFusion (Pfeiffer et al., 2021), are unsuitable due to oversimplified adapter aggregation. Our findings suggest the need for more targeted research on compositionality in scientific document summarization.
Figure 2. Illustration of explicit and implicit control attributes in CCSBench.
2. Related Work
2.1. Controllable Summarization
@@
Controllable summarization has attracted significant research interest, with efforts focused on guiding summaries using attributes such as length, topic, keywords, and sentiment. Early work explored the impact of oracle-provided attributes on summary quality (Fan et al., 2018; See et al., 2017). More recent studies have aimed at controlling individual attributes like length or keywords (Chan et al., 2021; He et al., 2020), but these methods often treat attributes independently, limiting their scalability and effectiveness in multi-attribute scenarios. MacSum (Zhang et al., 2023) represents a step forward by supporting multi-attribute control in the news and dialogue domains, where the content is relatively less technical. In contrast, scientific texts pose unique challenges for controllable summarization due to their high degree of linguistic precision and domain-specific complexity. CocoScisum (Ding et al., 2023) extends control to both length and keywords within scientific papers but remains constrained to explicitly defined attributes. Our work, CCSBench, advances the field by enabling fine-grained control over both explicit and implicit attributes, offering a broader and more nuanced evaluation framework tailored to scientific summarization.
2.2. LLMs for Scientific Document Processing
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of scientific document processing tasks, including idea generation (Si et al., 2025), literature review synthesis (Wang et al., 2024), research critique (Faizullah et al., 2024), and the production of scientific news articles (Liu et al., 2024b). These models also show promise in controllable text generation (Urlana et al., 2024; Zhou et al., 2024; Yun et al., 2025), though existing work primarily evaluates them under single-constraint settings (e.g., adjusting formality). Despite the growing need to tailor scientific content for diverse audiences, the ability of LLMs to perform compositional controllable generation – handling multiple constraints simultaneously – remains underexplored. To address this gap, we introduce the CCSBench benchmark, designed to systematically evaluate LLM performance under multi-attribute control. Our findings highlight the limitations of current models, particularly in handling implicit, conceptual attributes such as readability and aspect focus.
3. CCSBench Data Curation
3.1. Task Formulation
@@
Length specifies the desired number of words in the summary, ranging from concise to detailed, and catering to different audience needs based on their familiarity with the topic or time constraints. Length is divided into five bins, each representing a 50-word range (e.g., Bin 0: 0-50 words; Bin 1: 51-100 words).
Keywords ensures the inclusion of key terms from the source document, enabling readers to quickly grasp the core elements and assess the paper’s relevance. Keywords can be left empty or set as pre-defined terms to be included in the summary.
Readability adjusts syntactic complexity and vocabulary to match audience proficiency, making summaries more accessible or technical. This enhances comprehension and engagement. It has two levels: [normal, high].
Empirical Focus Level (Focus) adjusts emphasis between empirical results and theoretical contributions, with two levels: [low, high]. A high empirical focus highlights data-driven aspects like experiments and results, while a low empirical focus emphasizes theories, frameworks, and broader implications. This control tailors summaries to readers seeking practical insights or conceptual understanding.
3.2. Dataset
We construct CCSBench based on the arXiv dataset (Clement et al., 2019), a comprehensive collection of scientific documents from the widely used online preprint repository, arXiv. CCSBench focuses on papers in Computer Science and Artificial Intelligence. To ensure high-quality input with minimal noise, we use the introduction and conclusion sections (I+C) as the input text, a practice validated by prior research (Cachola et al., 2020; Meng et al., 2021) and our empirical analysis in Section 7. Following established practices in scientific summarization research (Cohan et al., 2018), we use abstracts as reference summaries, leveraging their widely recognized factuality and comprehensiveness (Stanford University, 2006).
The benchmark comprises four single-attribute control datasets and one compositional control dataset, each split into training, validation, and test sets using a 60/20/20 ratio. Dataset statistics are summarized in Table 1.
Table 1. CCSBench statistics.
@@ 3.2. Dataset
Table 2. Evaluation results validate the high quality of LLM-generated summaries in CCSBench.
Dataset FAC FLU CTRL ROUGE-1 ROUGE-2 ROUGE-L
Readability 0.98 0.93 0.92 53.88 21.81 41.31
Focus 0.91 0.93 0.93 51.25 34.23 41.41
3.3. Dataset Construction
@@
We begin with KWX (Fu, 2023), a dataset that provides keywords for each scientific document sourced from arXiv. From this, we select 3,700 papers in the subjects of Computer Science and Artificial Intelligence to form the foundation of our controllable summarization dataset, using their abstracts as natural reference summaries.
@@
3.4. Data Quality Validation
We assess the quality of GPT-4-generated summaries in CCSBench through automated and human evaluations. First, we compute ROUGE scores to assess fluency and content accuracy by comparing generated summaries with their original counterparts. Next, we conduct a human evaluation to further verify the summaries’ controllability and overall quality.
We sample 120 documents each from $D_{read}$ and $D_{focus}$ and recruit Amazon Mechanical Turk [^2] https://www.mturk.com/ annotators for binary assessments on three aspects: (1) Factual Consistency (FAC): whether the summary accurately reflects key points, (2) Fluency (FLU): whether it is natural and well-written, and (3) Controllability (CTRL): for Readability, whether the summary is easier to read than the reference; for Focus, whether empirically-focused summaries emphasize data collection, experimental setup, and results over conceptual aspects (details in Appendix B).
Table 2 shows high accuracy scores (¿0.9) across all metrics, confirming the quality of CCSBench ’s summaries. Rigorous annotator selection and strong inter-annotator agreement (¿0.85, Section 3.5) further validate these results.
3.5. Reliability of Human Evaluation
We implement rigorous measures to ensure the reliability of human evaluations. Annotators on Amazon Mechanical Turk are selected based on strict quality criteria, including a $\geq 98\%$ HIT approval rate and a minimum of 500 completed HITs, ensuring experienced and reliable contributors.
To validate annotation reliability, we conduct a pilot study with 20 randomly sampled instances for readability- and focus-controlled summaries. Each summary is evaluated by two independent annotators, and inter-annotator agreement is calculated. As shown in Table 3, all agreement scores exceed 0.85, confirming evaluation consistency. Based on these reliable pilot results, we scale the evaluation to two larger sets of 120 randomly sampled instances, each assessed by a single annotator, as shown in Section 3.4.
Table 3. Inter-annotator agreement scores for assessment questions on readability (Appendix B.1) and empirical focus (Appendix B.2).
Q1 Q2 Q3 Q4 Q5
Readability 1.00 0.85 0.95 0.90 -
Focus 0.95 1.00 0.85 0.90 0.95
4. Experimental Setup
4.1. Models
We evaluate a range of closed-source and open-source models for compositional controllable summarization, including GPT-3.5 (OpenAI, 2022), GPT-4 (OpenAI et al., 2024), GPT-4o (Hurst et al., 2024), LLaMA2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), and Flan-T5 (Chung et al., 2022). For closed-source models, we conduct few-shot experiments by randomly selecting three demonstrations per test sample. For open-source models, we explore parameter-efficient fine-tuning (PEFT) using Low-Rank Adaptation (LoRA)(Hu et al., 2021) to adapt them to our dataset. Additionally, we revisit controllable summarization baselines from the news domain, including hard prompt (HP)(Zhang et al., 2023) and soft prefix tuning (SP)(Li and Liang, 2021). Further implementation details, including training setups, are provided in Appendix A.
4.2. Evaluation Metrics
We evaluate all models based on both summarization quality and attribute controllability. We employ ROUGE (Lin, 2004) to evaluate the overall quality. To assess attribute controllability, we employ the following metrics tailored to each attribute.
@@ 4.2. Evaluation Metrics
We measure length controllability using the Mean Absolute Deviation (MAD)  (Liu et al., 2018), and the Pearson Correlation Coefficient (PCC)  (Liu et al., 2018), to evaluate the distance between the target and generated lengths.
@@
Keyword control is evaluated using the Success Rate (SR)  (He et al., 2020), which measures the fraction of specified keywords present in the generated summaries through exact matching after stemming.
@@
For readability, we calculate the Flesch-Kincaid Grade Level (FKGL)  (Kincaid et al., 1975), where a lower score indicates higher readability. The difference in FKGL scores ($\delta_{FKGL}$) between the high readability setting ($F_{high}$) and the normal readability setting ($F_{normal}$) reflects the model’s ability to control this attribute. A higher $\delta_{FKGL}$ is desirable.
@@
For empirical focus, we use GPT-4 to classify summaries into high or low empirical focus levels. Validated with 93% accuracy on a manually labeled dataset, GPT-4 predictions are used to compute the F1 score for each category, assessing the model’s focus control. Prompt details are provided in Appendix C.
Table 4. Performance comparison of representative language models on CCSBench, evaluating overall quality and four control attributes. Arch refers to model architectures, with ED for encoder-decoder and DO for decoder-only. Bold numbers indicate the best performance, while underlined numbers represent the second-best.
@@
5. Experiments